What is the radius of the path of an electron whose speed is 1x10^7 m/s in a magnetic field of magnitude 0.020 T when velocity is perpendicular to magnetic field?
1
Expert's answer
2012-11-12T11:44:59-0500
Lorentz force is equal to centrifugal acceleration * mass
mv^2/r = qvB
Hence
r = mv/(qB) = 9.1*10^(-31)*10^7/(1.6*10^(-19)*0.02) = 0.003м